# Modality Translation for Object Detection Adaptation Without Forgetting Prior Knowledge

### v1

@@ 5 Conclusion
In this supplementary material, we provide additional information to reproduce our work. The source code is provided alongside the supplementary material, and we are going to provide the official repository. This supplementary material is divided into the following sections: Detailed diagrams (Section 6), Quantitative Results (Section 7), in which we provide additional numerical results in terms of APs, and Qualitative Results (Section 8), in which we provide additional visualizations.
@@ 6 Detailed Diagrams
In Figure 6, we illustrate the proposed ModTr. This method involves using a single pre-trained model typically trained on the more prevalent data, i.e. RGB, by the input adaptation network. For clarity, the RGB modality (along with the RGB detector) is depicted in purple, while an adaptation block of ModTr is shown in green for one IR modality and in red for the other IR modality with different distribution.
Figure 6: Our proposed solution is based on using a single pre-trained model normally trained on the more abundant data (RGB) and then adapting the input through our ModTr block.
@@ 7 Quantitative Results
In this section, we provide further details for the different AP metrics of the experiments in the main manuscript. In Table 5, we compare the best ModTr with different image translation strategies. Some of the competitors use the RGB data for the translation, while others utilize bounding box annotations. Here, we can see that ModTr outperforms the competitors across different detectors and AP metrics, demonstrating its superiority in terms of classification and localization. For instance, when compared with HalluciDet using FCOS, which was the second best, our method shows an improvement of approximately $28$ better in terms of AP50, with a similar trend observed for RetinaNet on the LLVIP dataset. The gap is narrower with FasterRCNN, with an improvement of only around $2$ AP50. While the gap is smaller for the FLIR dataset, it remains consistent across all APs and detectors; for example, an improvement of $6$ AP50 for FasterRCNN, approximately $11$ AP50 for RetinaNet, and $11$ AP50 for FCOS. In terms of localization for the FLIR dataset, significant improvements are observed, such as around $7$ AP for FasterRCNN, $11$ AP for RetinaNet, and $11$ AP for FCOS when compared with HalluciDet. For the other methods, which do not rely on bounding boxes, substantial improvements are evident; for instance, our method exhibits an improvement of more than $11$ AP over FastCUT, with similar trends observed for other competitors.
Table 5: Comparison of the detection performance of three different detectors (FCOS, RetinaNet, and FasterRCNN) over different image-to-image methods to translate the infrared to RGB-like images and our proposed method (ModTr). The methods were evaluated on the infrared test set of LLVIP and FLIR datasets. The column RGB tells whether the method needs access to RGB images during training, and Box refers to the use of ground truth bounding boxes during training.
@@
FCOS RetinaNet FasterRCNN
@@
FCOS RetinaNet FasterRCNN
@@
Table 6: Comparison of the fine-tuning of the detector, the different non-parametric fusion strategies, and the fine-tuning version of the fusion. For each ModTr version, we have a correspondent fine-tuning train to update both the detector and ModTr network during training. The detection performance of three different detectors (FCOS, RetinaNet, and FasterRCNN) was evaluated on the infrared test set of LLVIP and FLIR datasets.
@@
FCOS RetinaNet FasterRCNN
@@
FCOS RetinaNet FasterRCNN
@@
ModTr 65.25 ± 0.33 32.24 ± 0.95 34.60 ± 0.38 64.96 ± 0.68 30.93 ± 0.50 33.85 ± 0.34 68.84 ± 0.40 34.77 ± 0.22 37.01 ± 0.15
In Table 7, we explore the potential of achieving comparable performance in detection task while significantly reducing the number of parameters by employing a smaller backbone for the translation network. The MobileNetv2 with only $6.6$ million additional parameters, achieves performance on par with ResNet34 which have $24.4$ million parameters. This reduction in parameters is even more pronounced when compared with a new detector on the desired modality. For example, a new FCOS detector would require more $33.2$ million parameters. This trend is similar over the various detectors and remains consistent for the FLIR dataset (Table 8).
Table 7: Study on the translation backbone describing the tradeoff between the number of parameters for the translation network and the detection performance for each detector on LLVIP dataset. The results comprise the three default detectors used (FCOS, RetinaNet, and FasterRCNN) in APs under different MobileNet and ResNet backbones for the translation network.
@@
Table 8: Study on the translation backbone describing the tradeoff between the number of parameters for the translation network and the detection performance for each detector on FLIR dataset. The results comprise the three default detectors used (FCOS, RetinaNet, and FasterRCNN) in APs under different MobileNet and ResNet backbones for the translation network.
@@ 8 Qualitative Results
In this section, we provide additional visualizations for all the methods, including images generated by our proposal and its detection results. First, in Figure 8, we present the detections in more detail, highlighting issues of those methods relying solely on translation, such as FastCUT, and some false positives when there is only FT. Subsequently, we provide additional visualizations for a batch of images processed by various detectors, i.e., in Figure 9 for FCOS, in Figure 10 for RetinaNet, and Figure 11 for FasterRCNN for both datasets.
@@
Figure 8: Bounding box predictions over different OD methods for infrared images on two benchmarks: LLVIP and FLIR. Yellow and red boxes show the ground truth and predicted detections, respectively. FastCUT [31] is an unsupervised image translation approach that takes as input infrared images (IR) and produces pseudo-RGB images. It does not focus on detection and requires both modalities for training. Fine-tuning is the standard approach to adapting the detector to the new modality. It requires only IR data but forgets the original knowledge of the pre-trained detector. Finally, ModTr, our approach focuses the translation on detection, requires only IR data, and does not forget the original knowledge so that it can be reused for other tasks.
@@
Figure 11: Illustration of a sequence of $8$ images of LLVIP and FLIR dataset for FasterRCNN. For each dataset, the first row is the RGB modality, followed by the IR modality, followed by FastCUT [31], and different representations created by ModTr and their variations.

### v3

@@ Acknowledgments
In this supplementary material, we provide additional information to reproduce our work. The code is available at https://github.com/heitorrapela/ModTr. This supplementary material is divided into the following sections: Detailed diagrams (Section 6), Different fusion strategies (Section 7), Quantitative Results (Section 8), in which we provide additional numerical results in terms of APs, Qualitative Results (Section 9), in which we provide additional visualizations and Additional Modality: Canny Edges (Section 10), in which we show experiments with edges extracted from IR images using Canny edge detector.
@@ 6 Detailed Diagrams
In Figure 6, we illustrate the proposed ModTr. This method involves using a single pre-trained detector model typically trained on the more prevalent data, i.e., RGB, and additional input adaptation network. For clarity, the RGB modality (along with the RGB detector) is depicted in purple, while an adaptation block of ModTr is shown in green for one IR modality and in red for the other IR modality with different distribution.
Figure 6: Our proposed solution is based on using a single pre-trained detector model normally trained on the more abundant data (RGB) and then adapting the input through our ModTr block.
@@
7 Different fusion strategies
Here, we present a comprehensive overview of the alternative fusion strategies we explored. In addition to the element-wise product fusion, which produced the best results for our study and was reported in the main manuscript, we also tested strategies ModTr+, and ModTr, detailed below:
ModTr+: The addition mechanism involves forwarding the input modality and summing it with the output of the translation network, as detailed in Equation 4. This residual connection of the input serves as regularization for the model, aiding the network in learning the missing information necessary for detector operation. This operator learns the new representation by amplifying pixel values when the weights of the translation representation tend toward $1$, or preserving the original information when they tend toward 00. Such a range of values is due to our modification on the U-Net [42] to generate, in which we changed the last layer to a Sigmoid function layer so we can better control the generated image to be closer to a real RGB-like image:
$\begin{split}\mathcal{L}_{\text{ModTr}_{+}}(x,Y;\vartheta)&=\mathcal{L}_{det}(f_{\theta}\left(h_{\vartheta}^{d}(x)+x\right),Y).\end{split}$ (4)
ModTr: The subsequent fusion mechanism draws inspiration from DenseFuse [28], which employs the relative importance of pixels as an attention mechanism for both the translation network’s output and input. This attention mechanism operates by providing a weighted average of the channels. The implementation details for such an operator can be found on [28]. Then, the proposed loss function is given by the following Equation 5:
$\begin{split}\mathcal{L}_{\text{ModTr}_{\oplus}}(x,Y;\vartheta)&=\mathcal{L}_{det}(f_{\theta}\left(\Phi(h_{\vartheta}^{d}(x),x)\right),Y),\end{split}$ (5)
with
$\Phi(x,\hat{x})=\frac{x\odot e^{x}+\hat{x}\odot e^{\hat{x}}}{e^{x}+e^{\hat{x}}}.$
@@ 8 Quantitative Results
In this section, we provide further details for the different AP metrics of the experiments in the main manuscript. In Table 5, we compare the best ModTr with different image translation strategies. Some of the competitors use RGB data for the translation, while others utilize bounding box annotations. Here, we can see that ModTr outperforms the competitors in terms of AP metrics across different detectors, demonstrating its superiority in terms of classification and localization. For instance, when compared with HalluciDet [31] using FCOS, which was the second best, our method shows an improvement of approximately $28$ better in terms of AP50, with a similar trend observed for RetinaNet on the LLVIP dataset. The gap is narrower with Faster R-CNN, with an improvement of only around $2$ AP50. While the gap is smaller for the FLIR dataset, it remains consistent across all APs and detectors; for example, an improvement of $6$ AP50 for Faster R-CNN, approximately $14$ AP50 for RetinaNet, and $11$ AP50 for FCOS. In terms of localization for the FLIR dataset, significant improvements are observed, such as around $7$ AP for Faster R-CNN, $11$ AP for RetinaNet, and $11$ AP for FCOS when compared with HalluciDet. For the other methods, which do not rely on bounding boxes, substantial improvements are evident; for instance, our method exhibits an improvement of more than $11$ AP over FastCUT [38], with similar trends observed for other competitors.
Table 5: Detection performance of ModTr versus baseline image-to-image methods to translate the IR to RGB-like images, using three different detectors (FCOS, RetinaNet, and Faster R-CNN). The methods were evaluated on IR test set of LLVIP and FLIR datasets. The RGB column indicates if the method required access to RGB images during training, and Box refers to the use of ground truth bboxes during training.
@@
FCOS RetinaNet Faster R-CNN
@@
FCOS RetinaNet Faster R-CNN
@@
Table 6: Detection performance (AP) of ModTr with different fusion strategies versus baseline fine-tuning (FT) of the detector, using three different detectors (FCOS, RetinaNet, and Faster R-CNN. The methods were evaluated on IR test set of LLVIP and FLIR datasets.
@@
FCOS RetinaNet Faster R-CNN
@@
FCOS RetinaNet Faster R-CNN
@@
ModTr 65.25 ± 0.33 32.24 ± 0.95 35.49 ± 0.94 64.96 ± 0.68 30.93 ± 0.50 34.27 ± 0.27 68.84 ± 0.40 34.77 ± 0.22 37.21 ± 0.46
In Table 7, we explore the potential of achieving comparable performance in detection tasks while significantly reducing the number of parameters by employing a smaller backbone for the translation network. The MobileNetv2 with only $6.6$ million additional parameters, achieves performance on par with ResNet34 which have $24.4$ million parameters. This reduction in parameters is even more pronounced when compared with a new detector on the desired modality. For example, a new FCOS detector would require more $33.2$ million parameters. This trend is similar over the various detectors and remains consistent for the FLIR dataset (Table 8).
Table 7: Detection performance of ModTr with different backbones for the translation networks with different numbers of parameters, using three different detectors (FCOS, RetinaNet, and Faster R-CNN). The methods were evaluated on IR test set of LLVIP dataset.
@@
Table 8: Detection performance (AP) of ModTr with different backbones for the translation networks with different numbers of parameters, using three different detectors (FCOS, RetinaNet, and Faster R-CNN). The methods were evaluated on IR test set of FLIR dataset.
@@ 9 Qualitative Results
In this section, we provide additional visualizations for all the methods, including images generated by our proposal and its detection results. First, in Figure 8, we present the detections in more detail, highlighting issues of those methods relying solely on translation, such as FastCUT, and some false positives when there is only FT. Subsequently, we provide additional visualizations for a batch of images processed by various detectors, i.e., in Figure 9 for FCOS, in Figure 10 for RetinaNet, and Figure 11 for Faster R-CNN for both datasets.
@@
Figure 8: Bounding box predictions over different OD methods for infrared images on two benchmarks: LLVIP and FLIR. Yellow and red boxes show the ground truth and predicted detections, respectively. FastCUT [38] is an unsupervised image translation approach that takes as input infrared images (IR) and produces pseudo-RGB images. It does not focus on detection and requires both modalities for training. Fine-tuning is the standard approach to adapting the detector to the new modality. It requires only IR data but forgets the original knowledge of the pre-trained detector. Finally, ModTr, our approach focuses on the translation on detection, requires only IR data, and does not forget the original knowledge so that it can be reused for other tasks.
@@
Figure 11: Illustration of a sequence of $8$ images of LLVIP and FLIR dataset for Faster R-CNN. For each dataset, the first row is the RGB modality, followed by the IR modality, followed by FastCUT [38], and different representations created by ModTr and their variations.
10 Additional Modality: Canny Edges
As described in [1], edges extracted with a Canny Edge detector from IR images can be used as an additional modality. Thus, in our work, we provide some qualitative results with it as well, as illustrated by Figure 12 and quantitative results in Table 9.
Table 9: Detection performance of Faster R-CNN Zero-Shot, Fine-Tuning, and ModTr on Edge Modality for LLVIP and FLIR Datasets. Edges from IR images were extracted using the Canny Edge Detector.
Method Test Set Canny Edges (Dataset: LLVIP)
Faster R-CNN
[…]
Zero-Shot 14.33 ± 0.00 08.15 ± 0.00 08.01 ± 0.00
Fine-Tuning 87.39 ± 1.74 55.50 ± 2.61 54.48 ± 2.40
ModTr 81.23 ± 1.53 48.66 ± 1.57 47.06 ± 1.18
Method Test Set Canny Edges (Dataset: FLIR)
Faster R-CNN
[…]
Zero-Shot 18.43 ± 0.00 06.68 ± 0.00 08.44 ± 0.00
Fine-Tuning 60.43 ± 2.05 26.38 ± 0.28 30.88 ± 0.88
ModTr 51.81 ± 1.41 22.14 ± 0.83 25.59 ± 0.83
Detector: Faster R-CNN
LLVIP Test Dataset (Canny Edge Modality)
FT Refer to caption
ModTr Refer to caption
FLIR Test Dataset (Canny Edge Modality)
FT Refer to caption
ModTr Refer to caption
Figure 12: Illustration of a sequence of $8$ images from LLVIP and FLIR Canny edges test set for Faster R-CNN. The first row is the FT on the Canny edges modality, and the second row is the result of ModTr for the LLVIP. The third row is FT on the Canny edges for FLIR, and the fourth row is the result of ModTr.
Here, we clarify that our approach is dependent on the zero-shot capability of the model to incorporate knowledge of the translation network. So, we can incorporate other modalities as well as the IR, but if the new modality does not have a good zero-shot like the Canny Edges, our method can incorporate the new knowledge but not surpass the fully fine-tuning (FT). But it is also important to mention that even without surpassing for edges, we were able to keep prior knowledge and reach good performance even though we still have a gap for the edges version when compared with FT.